OpenStudy (anonymous):

find the area of a right triangle with given leg l and the hypotenuse h. round decimal answers to the nearest tenth 1. l=8 m, h= 16 m 2. l = 9 yd, h= 12 yd 3. l= 21in., h + 29 in. some one help me please!!!

OpenStudy (anonymous):

1) a^2 + b^2 =c^2 a and b are the legs and c is the hypotenuse. 8^2 +x^2 = 16^2 64+ x^2 = 256 x^2 = 192 x= 13.9

OpenStudy (anonymous):

thanks can you show me the other 2 please

OpenStudy (anonymous):

yeah, just so you know that is just the length for the other leg, after you find that length you have to plug all the numbers in to the formula for finding the area of a right triangle. 2) a^2 + b^2 = c^2 9^2 + x^2 = 12^2 81+ x^2 = 144 x^2 = 63 x = 7.9 3) a^2 + b^2 = c^2 21^2 + x^2 = 29^2 441 + x^2 = 841 x^2 =400 x=20

OpenStudy (anonymous):

thanks can u hehelp me with another few problems  and it says find the area of the figure. round decimal answers to the nearest tenth

OpenStudy (anonymous):